### v1

# When Vision Fails: Text Attacks Against ViT and OCR

Abstract
While text-based machine learning models that operate on visual inputs of rendered text have become robust against a wide range of existing attacks, we show that they are still vulnerable to visual adversarial examples encoded as text. We use the Unicode functionality of combining diacritical marks to manipulate encoded text so that small visual perturbations appear when the text is rendered. We show how a genetic algorithm can be used to generate visual adversarial examples in a black-box setting, and conduct a user study to establish that the model-fooling adversarial examples do not affect human comprehension. We demonstrate the effectiveness of these attacks in the real world by creating adversarial examples against production models published by Facebook, Microsoft, IBM, and Google.
1 Introduction
Advances in adversarial examples [1, 2, 3] for text-based machine learning systems have led to imperceptible perturbation techniques targeting the encoding of text [4, 5]. These attacks leverage uncommon Unicode encodings to subvert text models without visual indication to users [6]. Existing defenses against encoding attacks achieve protection by unifying the visual and encoding pipelines; specifically, the Vision Transformers (ViTs) architecture can be used to build robust new models [7], and Optical Character Recognition (OCR) can be used to retrofit existing models with defenses [4]. Such defenses seek to ensure that text inputs which are visually-identical when rendered produce equivalent outputs in defended models.
We show that these defenses are inadequate. We propose a technique to encode text perturbations in a way that, once rendered, the image of the rendered text will contain adversarial perturbations that defeat visual defenses. These adversarial examples operate in the visual text domain, meaning that visual inputs are generated exclusively from rendered text such that it is impossible to perturb arbitrary pixel values. However, by leveraging combining marks from the Unicode specification, we can craft small, targeted visual perturbations that appear on the rendered image of text. While these marks are not sufficiently visually significant to impact a human reader’s understanding of text, the manipulated pixels in the image domain of rendered text enable targeted attacks on model outputs.
Text attacks on ViT and OCR have broad implications. For example, we demonstrate extensions to attacks on toxic content detection [8] and machine translation [9] that succeed despite ViT and OCR defenses against encoding attacks. The techniques we propose can even be used to affect the physical domain such as disclosures submitted within legal cases. If the disclosing legal team adds appropriate text perturbations to printed disclosure documents, the text can form visual adversarial examples that are incorrectly processed by OCR systems despite being readable by humans. This would prevent searchability thus impeding the opposing legal team in the discovery process.
In this paper we make the following contributions:
  • We introduce a novel form of adversarial example against machine-learning models that process rendered text, including Vision Transformers and Optical Character Recognition models, which are encoded in the textual domain but operate in the visual domain.
  • We demonstrate that previously-proposed defenses [4, 7, 10] for encoding attacks against text models including ViT, OCR, and neural encoders are insufficient.
  • We conduct a user study to validate that our adversarial examples do not impact human readability.
  • We propose definitive defenses for diacritical attacks against Unicode-based visual models.
2 Background
Figure 1: The visualization gap, against which visual models claim to be robust, continues to be source of adversarial examples via diacritics as in this toxic content classifier example.
The Visualization Gap. Traditionally, machine learning models processing text such as natural language operate directly upon some encoding of the input text. This may take the form of input embeddings as vectors representing words, characters, or learned subword components created by parsing Unicode inputs [6]. However, unlike models, humans do not directly consume encoded text. Rather, text is rendered and then visually presented to human users.
It is here that a security design flaw arises: the relationship between encoded text and rendered text is not bijective. That is, a visual rendering could be represented by many unique text encodings. Formally,
$\forall{T}\in t,\ \ U(t)\nLeftrightarrow\{v(t)\}$ (1)
where $T$ is the set of all possible text sequences, $U$ is the function generating the set of all possible Unicode representations of a text, and $v$ is the visual rendering of a text.
Consider, for example, the presence of invisible characters such as Unicode’s Zero-Width Space (ZWSP); these characters have no effect on the rendering of most text and yet change the encoded representation. Visually-identical characters, known as homoglyphs, can also be used interchangeably, and control characters can be used to delete and reorder characters.
The difference between the encoding and visualization of text can be used to create adversarial examples against models that operate directly upon some form of textual input [4, 5], improving the stealth of earlier techniques leveraging misspelling or paraphrasing [11, 12, 13, 14]. The visualization gap is depicted in Figure 1.
Defense Through Vision. To defend against adversarial examples that exploit the visualization gap, model designers must seek to unify the text encoding and visualization pipelines. That is, designers must seek to build or augment models such that:
$\forall{T}\in t,\ \ E(U(t))=\{t^{\prime}\}\Leftrightarrow\{v(t)\}$ (2)
where $E$ generates the set of embeddings for the encoded values taken as input.
One simple but effective way to accomplish this on existing models is to render text inputs and process the resulting images through OCR as a pre-processing step prior to model inference [4]. In effect, this provides an automated system that maps fixed visual renderings to a common encoded input. The inference pipeline in this setting is: encoded input $\rightarrow$ rendered image $\rightarrow$ text $\rightarrow$ model.
For greenfield models, Vision Transformers may be the preferred defense approach as no compute-intensive pre-processing model is required. ViTs operate upon images as input, and operating directly upon rendered images as embeddings yields both good performance and defense-by-design against attacks exploiting the visualization gap [7]. The inference pipeline in this setting is: encoded input $\rightarrow$ rendered image $\rightarrow$ model.
Finally, neural encoders offer new NLP models some robustness against Unicode perturbations [10]. Although they do not operate in the visual domain, neural encoders are a form of learned embedding that map relationships between Unicode characters such that encoded values that would look similar if rendered should result in similar embeddings. The inference pipeline in this setting is: encoded input $\rightarrow$ neural embedding $\rightarrow$ model.
Figure 2: Our attack: adversarial examples in the visual domain encoded in the textual domain.
3 Methods
Exploiting Diacritics. In the image domain, adversarial examples are typically crafted by slightly perturbing the values of key pixels often identified through a gradient-based approach [3]. While such an approach would in theory work against ViTs and OCR models, the visual text domain has the added constraint that the input images are generated through rendering text. Therefore, it is not possible to arbitrarily perturb pixel values, as pixel values are not directly encoded; rather, the pixels are generated from encoded text.
Diacritics – also known as diacritical marks or accents – are small marks that can be placed on top of other written letters. These marks serve different purposes across different linguistic families, but often serve to modify the pronunciation or meaning of words. Examples of diacritics include the Spanish ñ, the German ä, and the French è. Diacritics are most common in European languages, but are also used to aid in the romanization of non-Latin scripts such as the case of pinyin for Mandarin.
In Unicode, characters commonly used with diacritical marks typically have a dedicated character – or code point – in the Unicode specification. However, the concept of combining diacritical marks also exists. These characters modify the character immediately preceding them to add the specified diacritical mark. The Unicode specification defines 256 such combining characters [15, 16, 17, 18, 19].
Combining diacritical marks provide a method by which arbitrary characters can be subtly perturbed in the image domain. These marks will produce a noticeable visual effect, but as we will later demonstrate through a user study these marks do not affect a human’s ability to read the text. However, just as changing particular pixel values can cause general image classifiers to fail, so too can the perturbed pixel values created by diacritical marks affect ViTs and OCR models. Although the size of the perturbation space is smaller than the more continuous space provided by color pixels in unconstrained images, we will later demonstrate through a series of experiments that this space is sufficiently large to generate adversarial examples for text. This attack is visualized for ViT and OCR pipelines in Figure 2.
  Input: text $\mathbf{x}$, reference output $\mathbf{y}$, model $\mathcal{M}$, perturbation budget $\beta$, diacritics list $\mathbf{D}$
[…]
  Result: Encoded text visually similar to x which is an adversarial example against $\mathcal{M}$ when rendered
  procedure perturb ($p$)
   $\bar{x}:=\mathbf{x}$
   for $n:=0$ to $\beta$ do
    $d,i:=p_{n}$
    if round($i$) $\geq 0$ then
     $\bar{x}=$ insert($\bar{x},\mathbf{D}_{d},$round($i$))
    end if
   end for
   return $\bar{x}$
  end procedure
Randomly initialize   $\mathbf{P}:=\{\mathbf{p_{0}},\dots,\mathbf{p_{s}}\}$ ,
where $\mathbf{p_{n}}:=[(d_{0},i_{0}),...,(d_{\beta},i_{\beta})]$ ,
where $d_{n}\sim\mathcal{U}(0,|\mathbf{D}|),\ i_{n}\sim\mathcal{U}(-1,|\mathbf{x}|)$
$\rhd\ \mathcal{U}$ is uniform dist.
  if $\mathcal{M}$ is classifier then
[…]
  else
     $\mathcal{F}(\hat{x})=\mathcal{S}(\mathcal{M}(\mathbf{\hat{x}}),\mathbf{y})$
  end if
  for $i:=0$ to $m$ do
     for $j:=0$ to $s$ do
[…]
        $R\sim\mathcal{U}(0,|\mathbf{p_{j}}|)$
        $\mathbf{\hat{p}_{j}}:=\mathbf{p_{j}}$
        for $k:=0$ to $|\mathbf{p_{j}}|$ do
           $r_{j}\sim\mathcal{U}(0,1)$
[…]
              $\mathbf{\hat{p}_{j_{k}}}=\mathbf{p_{a_{k}}}+F\times(\mathbf{p_{b_{k}}}-\mathbf{p_{b_{k}}})$
           end if
        end for
        if $\mathcal{F}(\mathbf{\textsc{perturb}(\hat{p}_{j}}))<\mathcal{F}(\textsc{perturb}(\mathbf{p_{j}}))$ then
           $\mathbf{p_{j}}=\mathbf{\hat{p}_{j}}$
        end if
     end for
  end for
  $\mathbf{\bar{f}}:=\{\mathcal{F}(\textsc{perturb}(\mathbf{p_{0}})),$
           $\dots,\mathcal{F}(\textsc{perturb}(\mathbf{p_{s}}))\}$
  return  $\textsc{perturb}(\mathbf{P}_{\text{argmax}(\mathbf{\bar{f}})})$
Algorithm 1 Black-box generation of visual text adversarial example
Attack Technique. Visual adversarial examples encoded as text can be generated in a black-box model using a gradient-free optimization technique. We built our technique using differential evolution, a genetic optimization algorithm [20]. Adopting a similar approach previously proposed for imperceptible perturbations [4, 21], we use differential evolution to minimize similarity with the reference output for sequence-to-sequence tasks, and to minimize the output probability of the target class for classification tasks.
The algorithm to generate visual adversarial examples encoded as text is provided as Algorithm 1. In summary, this algorithm takes as input a string to perturb, the target output class or reference output sequence, a visual text model, a perturbation budget representing the maximum number of injected diacritics, and a set of diacritics from which to select. A set of optimization parameters specific to differential evolution are also required by the algorithm [20], although reasonable defaults may be used across repeated invocation. In the case of classification models, this algorithm will return a version of the input string perturbed with diacritics up to the allowed budget that minimize the output probability of the supplied target class. In the case of sequence-to-sequence models – such as the machine translation task – the algorithm minimizes the supplied similarity metric with the supplied reference output.
Since the adversarial example generation algorithm does not rely on gradients, it works in a black-box setting, i.e. the attack does not require access to model weights. Consequently, this algorithm can be used both for local models and on commercial models via APIs as demonstrated in the following section.
4 Evaluation
The experimental evaluation of visual text adversarial examples requires examining two claims: first, that the adversarial examples effectively degrade the performance of a broad set of models, and second, that the adversarial examples do not affect human comprehension.
To examine performance against models, we will first generate adversarial example for an OCR model and measure performance degradation. We will then place this model into a pipeline that renders text input, performs OCR, and then calls the downstream model for both machine translation and toxic content detection and evaluate the performance of adversarial examples. Next, we will craft and measure adversarial examples for a ViT performing machine translation directly on rendered text. Finally, we will demonstrate an extension of these attacks to a neural encoder model performing question answering. A summary of the experimental results is given in Table 1. These results are presented over varying budgets in Figure 4, with additional metric-specific visualizations given in the Appendix.
To examine the effect on human comprehension, we will present a user study in which participants are evaluated for their understanding of the adversarial examples produced in the previous experiments.
All adversarial examples were generated on a cluster of machines each equipped with a Tesla P100 GPU and Intel Xeon Silver 4110 CPU running Ubuntu. We followed Algorithm 1 for example generation, selecting a a population size of 32, 10 iterations, a crossover probability of 0.7, a differential weight dithering from 0.5 to 1, and a varying budget ranging from 0 to 5. For each experiment, we generate 500 adversarial examples for each budget value. Perturbations are generated from the subset of diacritical marks in Unicode supported by the Microsoft Arial Unicode font [22], which is U+300-U+346 and U+360-U+361. All experimental code and results are made available at github.com/nickboucher/diacritics.
Model Baseline Perf Adv. Ex. Perf $\beta=5$ Attack Perf Drop TrOCR [23] 0 12 12 TrOCR-FairSeq [24] 60.9 24.2 60.3% TrOCR-IBM Toxic [25] 87.6 31.8 63.7% ViT FairSeq [7] 57.9 24.5 57.7% CANINE SQuAD [10] 67.6 5.0 92.6% Table 1: Model performance against adversarial examples of diacritical mark budget 5.
Figure 3: Evaluation of diacritic adversarial examples against TrOCR.
TrOCR. TrOCR is a transformer model [26] published by Microsoft implementing OCR [23]. TrOCR achieves state-of-the art performance on text recognition and leverages the modern transformer architecture to implement image understanding and text generation in a single model.
Our first experiment evaluates the performance of TrOCR against diacritics injected via Algorithm 1. For each example, we render the input using the Microsoft Arial Unicode font [22] and pass the rendered image to the TrOCR model for inference. In this experiment, we use negative Levenshtein distance of the model output with the adversarial input as the similarity metric. If the TrOCR model is highly performant, we would expect a trivially small average distance between the input and output when the attack budget is zero (representing no diacritic injections); indeed, this is what we see in the results. The experiments seek to generate adversarial example using diacritics injected into inputs sampled from the grammatically-correct validation split of the CoLA dataset [27], selected as a reproducible sample of short English-language inputs.
In Figure 3 we plot the distance between the output and both the perturbed and unperturbed input. Both measures grow with the budget, indicating that the model neither removes diacritical marks nor recognizes them as diacritics.
[…]
In Figure 6 we plot the average chrF score [29] – a measure of translation quality selected as the similarity matric for this experiment – relative to the reference translation as computed by sacreBLEU [30]. The results indicate a decreasing translation quality with increasing diacritics budget.
Defended Toxic Classifier. Similarly, we evaluated another model defended from Unicode attacks via TrOCR. In this experiment, the downstream model is IBM’s toxic content classifier [25]. We use the Wikipedia Detox Dataset [31] as the inputs from which we craft adversarial examples.
The classification performance of the defended model over varying attack budgets is shown in Figure 6. There is, again, a negative relationship between the performance of the model and the attack budget. In fact, with a budget of just two diacritical marks, the performance of the model falls below random (i.e. 50% correct classification).
[…]
In Figure 8 we plot the average chrF score – again chosen as the similarity metric – over the perturbation budget. Similar to the TrOCR-defended machine translation experiment, we see a significant negative correlation between translation quality and diacritics perturbation budget.
CANINE Question Answering. Neural encoders are a recent NLP tool designed to replace dictionaries with embeddings learned directly from Unicode code points. Although these models do not unify the visual and encoded pipelines for text input, the learned embeddings claim to offer better flexibility towards otherwise unknown characters. In this experiment, we evaluate whether diacritical mark attacks can subvert models leveraging these neural encoders.
We evaluated a question answering model leveraging the CANINE neural encoder proposed by Google [10]. Specifically, we evaluated a transformer model [33] trained on the SQuAD dataset [34]. Consistent with benchmarks against this dataset, we select F1 score as the similarity metric. The results, shown in Figure 8, show a sharp decline in F1 score with increasing diacritical perturbations.
Figure 4: Performance of evaluated models across perturbation budgets relative to no perturbations.
5 User Study
Having shown that OCR and ViT models perform poorly when interpreting text containing visual perturbations, and that these failures can be used to craft adversarial examples for downstream tasks, we conduct a user survey to understand how human comprehension is affected by diacritical perturbations.
User Study Methodology. The survey contained two main questions, each with multiple sub-components. The questions are included as Figures 10 and 11.
First, to measure human ability to read text containing diacritics, we ask respondents to retype sentences containing diacritics without using any diacritics or other marks. We selected five short, benign sentences as examples to measure comprehension.
Second, to measure human comprehension of text containing diacritics, we asked respondents to identify whether a short sentence is considered toxic using the toxicity definition of Google’s Perspective API [35]. We hand-selected examples to choose short sentences that contained toxic examples with comparatively milder language. We used 12 example sentences in total, evenly split between six toxic and six non-toxic examples. The sentences were randomly ordered within the survey question to avoid biasing respondents.
We surveyed 200 people using Amazon Mechanical Turk, limiting the survey to include only MTurk workers with at least a U.S. high school education since our questions concerned critical interpretation of text. The survey was advertised as “Answer a short (estimated 3-4 min) survey about reading and interpreting short sentences”, intentionally avoiding any mention of diacritics, accents, or unusual marks. We identify six identical low-quality responses that failed quality-control screening, leaving a total of $n=194$ responses.
Example
Ground Truth
Toxicity Label
Model Label
Human Reader
Label
Human Reader
Performance
Model
Correct
Humans
Correct
[Uncaptioned image] True False True 93.8%
[Uncaptioned image] True False True 94.8%
[Uncaptioned image] True False True 96.4%
[Uncaptioned image] False True False 99.0%
[Uncaptioned image] False True False 99.0%
[Uncaptioned image] False True False 99.0%
[Uncaptioned image] True True True 80.4%
[Uncaptioned image] True True True 97.9%
[Uncaptioned image] True True True 93.3%
False False False 98.5%
[Uncaptioned image] False False False 99.5%
[Uncaptioned image] False False False 99.0%
Table 2: Comparison of OCR and ViT model performance and human performance detecting toxicity in adversarial examples. A ‘True’ label means the adversarial example was identified as toxic, while ‘False’ means the example was considered non-toxic.
We paid respondents \$1.50 USD to complete the survey, which had an average completion time of just under three minutes and maximum completion time of up to ten minutes. This amount was chosen to reflect a fair hourly wage.
User Study Results. We conclude that diacritical marks do not significantly affect a human user’s ability to comprehend the adversarial examples generated in the prior experiments. User responses were highly accurate in both reading and interpreting text containing diacritical marks, indicating that OCR and ViT models struggle with a text interpretation task humans find simple.
We find that the overwhelming majority of human readers are able to correctly read and repeat the example sentences. Respondent accuracy rates for the three example sentences ranged from 93.3% to 97.9% . In evaluating what constitutes an ‘accurate’ response, we ignore minor punctuation differences that seem unlikely to have been caused by diacritical marks (e.g., lack of capitalization, quotation marks around the sentence, a lack of a period at the end of the sentence, etc.).
Inaccurate responses mostly substituted a word that contained a diacritical mark for a similar but incorrect word (e.g., “tube” instead of “tub”, or “John laid” instead of “John lay”). While a small minority of respondents struggled to identify certain words containing diacritics, we conclude that the vast majority of human readers are capable of comprehending text regardless of diacritical marks.
Table 2 contains the twelve examples used in the survey, evenly divided between six examples the TrOCR-defended toxic content classifier labeled correctly (based on the ground-truth toxicity label from the Wikipedia Detox Dataset [31]) and six it did not. Human accuracy was maintained through all adversarial examples suggesting that human accuracy is unaffected by the presence of diacritical marks, even as OCR-defended model accuracy drops.
6 Discussion
Ethics. All attacks performed were run against local models and did not target any online services. We leveraged laboratory GPU resources for attack generation and did not use cloud services. Our user study was reviewed and approved by our Institutional Review Board, and no personal information was tracked from survey participants. We compensated users via Amazon Mechanical Turk at fair market prices. Since the study involved viewing toxic content, per Mechanical Turk policy we included a warning in the survey title that the assignment contains “language that may be considered rude or disrespectful.”
Limitations. The adversarial examples in this work produce minor visual artifacts by design. These examples contrast the assumptions made by Boucher et al. [4] in which no visual artifacts were permitted in encoding attacks. However, the results of our user study indicate that the diacritic perturbations produced do not affect the ability of humans reading text, thus giving motivation to break the imperceptibility assumption of encoding attacks.
Defenses. The attacks described in this paper all use Unicode diacritical marks to encode visual perturbations in the textual domain. Diacritics carry important linguistic value – particularly for pronunciation purposes – and cannot be disallowed in inputs without breaking internationalization.
[…]
In settings where diacritical marks can’t be removed without hurting the performance of the model, the next best option is to perform adversarial training on the ViT, OCR, or neural encoder processing textual inputs. Yet, this will not come without additional risks via invariance-based adversarial examples where the model will learn to detect characters that are no longer human decipherable [36].
Rendering Design. In the visual text domain, setting-specific engineering details become pseudo-hyperparameters of the model. In this setting, it is no longer sufficient to communicate the model architecture and weights, but rather model providers must provide a thorough implementation or explanation of text rendering, chunking, and canvas handling along with their trained models.
One engineering challenge introduced by visual models is the need to implement text rendering onto fixed-sized canvases. By the nature of machine learning in the visual domain, model inputs must take the form of a single, fixed size image as specified by the model. This image size, which we will call the canvas dimension, introduces a key hyperparameter of the model that must be optimized. Additionally, one must also specify the font and font size that will be used for rendering. The implementation must also account for how to fit text of variable length onto the canvas, which may be further complicated if the font is not fixed width.
When there is too much text to fit onto the canvas, the canvas cannot be resized to make it fit; this would necessitate the image being further resized before inference, and the text becoming too small for the model to optimally recognize. Instead, long text inputs must be chunked into appropriately sized sections that each fit on their own canvas. These inputs must be processed separately and combined after inference (a process which must account for potential false whitespace caused by the edge of the canvas). Further, if words are broken across canvases the model will likely lose the performance benefits of recognizing adjacent characters of higher likelihood. We provide a visualization of this in Figure 9.
In addition to the aforementioned challenges, the text rendering engine selected must also provide robust rendering functionality for chosen encoding – such as Unicode – inclusive of support for control characters, bidirectional text, and any other details of the text specification.
7 Related Work
Text-Based Adversarial Examples. While most adversarial examples proposed in prior work have been image classification attacks [1, 3, 37], recent work has demonstrated the existence of adversarial examples against text-based ML models. In particular, the accuracy of text classification models decreases on noisy inputs and it is possible to deliberately craft adversarial texts targeting systems reliant on text understanding [11, 12, 13, 14, 38, 39].
In contrast to prior work presenting visibly perceptible perturbations, Boucher et al. [4] previously showed that the discrepancy between text encoding and text visualization can be exploited as an attack vector against text-based machine learning models. They proposed optical character recognition (OCR) models as a “catch-all defense” against such attacks, arguing that the visual nature of the input would mitigate differences in visual and text encoding representation that left text-based processing systems vulnerable.
While Boucher et al. evaluated OCR as a defense against imperceptible perturbations, diacritics present a new class of attacks that are often scarcely more perceptible than invisible perturbations but have a substantial impact on OCR accuracy. Salesky et al. [7] showed that Arabic diacritics and other minor text modifications substantially degrade the performance of text-based models, but concluded that ViTs achieve “relatively robust” performance against diacritized text. In this paper we demonstrate that both OCR systems and ViTs are indeed vulnerable to the Unicode-based adversarial examples introduced by Boucher et al., and that such examples significantly impact the performance of these models.
Attacks on OCR. Most prior work in breaking OCR systems has focused on visual CAPTCHAs [40, 41], but OCR is commonly used in a variety of preprocessing tasks. Recently, Chen et al. [42] demonstrated a class of attacks on OCR using adversarial watermarks in a twist on classic adversarial image examples. The commonality of watermarks on inputs to OCR systems makes this attack effectively imperceptible to humans, though this attack relies on non-text image data in inputs. In contrast, in this work we present a class of attacks in which the adversarial examples are encoded directly into text rather than perturbing an image’s background. To the best of our knowledge, this is the first such attack that manipulates pre-rendered text directly, using the resulting image as an adversarial example.
Toxic Content Detection. Identifying toxic content online has proven to be one of the most pervasive yet elusive research questions in natural language processing. A rapidly growing area of research, toxic content detection is challenging even when classifying standard Unicode characters in the English language, and toxic content classifiers are generally not robust against adversarial examples [43, 44]. Users online commonly deploy a variety of simple text modifications to bypass platform filters: Kurita et al. [43] gives the adversarial example of writing “s*ut up” in the place of ‘shut up’. While most adversarial examples against toxic content detection are visually perceptible, Boucher et al. [4] showed it is also possible to modify the Unicode text encoding of toxic content in a way that is visually imperceptible but fools content detection models. In this paper, we demonstrate that text containing diacritical marks presents a further class of attacks against toxic content detection systems.
8 Conclusion
We have presented a novel method of attacking language models operating in the visual domain by encoding adversarial examples in the textual domain. Once rendered, these examples will generate small visual artifacts that drastically decrease model performance. We accomplish this through the injection of Unicode’s combining diacritical marks via a gradient-free optimization designed to minimize target model performance over the perturbed input. We generate such adversarial examples for real-world models published by Microsoft, Facebook, IBM, and Google finding the attacks to significantly degrade model performance. We also conduct a user study which concludes that the comprehension of human users is not affected by the diacritic injections used in our attacks. In order to defend against such attacks, we conclude that model designers should remove combining diacritical marks from inputs prior to inference.
Appendix
Figure 5: Evaluation of diacritic adversarial examples against FairSeq’s EN->FR translation model defended by TrOCR.
Figure 5: Evaluation of diacritic adversarial examples against FairSeq’s EN->FR translation model defended by TrOCR.
Figure 6: Evaluation of diacritic adversarial examples against IBM’s toxic content detection model defended by TrOCR.
Figure 7: Evaluation of diacritic adversarial examples against the Visual Representation fork of FairSeq for DE->EN translation.
Figure 7: Evaluation of diacritic adversarial examples against the Visual Representation fork of FairSeq for DE->EN translation.
Figure 8: Evaluation of diacritic adversarial examples against SQuAD transformer using Google’s CANINE neural encoder.
Figure 9: An example of three chunking options to render the text “uncharted backwaters of the” onto a series of fixed-size canvases for inference. Option A breaks letters across canvases, which will cause failures during inference. Option B breaks the word backwaters across more canvases than necessary, decreasing the model’s ability to recognize likely adjacencies. Option C is likely the strongest choice in this example, but as with all other options the model pipeline must account for whitespace added by chunking after inference.
Figure 10: First survey question measuring human readers’ comprehension of text containing diacritics.
Figure 11: Second survey question measuring human readers’ ability to identify toxic text containing diacritics.